OpenStudy (anonymous):

Solve the following equation: 6x2 + 23x = –20.

OpenStudy (anonymous):

i got x=1/3 and x=-14

OpenStudy (anonymous):

it looks like this\[6x ^{2}+23x=-20\]

OpenStudy (anonymous):

since ques. is of form ax^2+bx+c=0,Find D=Discriminant=b^2-4ac,then find x=(-b+(D^1/2))/2a OR x=(-b-(D^1/2))/2a.

OpenStudy (anonymous):

ok i will try it